A body excutes SHM whose period is 16s.Two second after it passes the equilibrium position,its velocity is 1m/s.The amplitude of SHM is.

Grade:12

2 Answers

Anil Pannikar AskiitiansExpert-IITB
13 years ago

Dear Sunny,

 

 as y = Asinwt

so, dy/dt = v = Awcoswt

here w = 2Π/T = Π/8 , v =1and t = 2

plug in all values

 

Please feel free to ask your queries here. We are all IITians and here to help you in your IIT JEE preparation.

All the best.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar  respectively : Click here to download the toolbar..

 

Askiitians Expert

Anil Pannikar

IIT Bombay

 

 

vikas askiitian expert
13 years ago

 x=Asinwt                                                                           

 thus v=dx/dt =Awcoswt                                                      (A is amplitude)

 w is angular frequency =2pie/T =pie/8                                   (here T is time period)

                               w=pie/8

 after 2 seconds its velocity is 1

                         1 =Awcoswt =awcospie/4

                         A=toot2/w=8root2/pie is required ans